# Minimax rates in variance and covariance changepoint testing

### v1

@@ Abstract
We study the detection of a change in the spatial covariance matrix of $n$ independent sub-Gaussian random variables of dimension $p$. Our first contribution is to show that $\log\log(8n)$ is the exact minimax testing rate for a change in variance when $p=1$, thereby giving a complete characterization of the problem for univariate data. Our second contribution is to derive a lower bound on the minimax testing rate under the operator norm, taking a certain notion of sparsity into account. In the low- to moderate-dimensional region of the parameter space, we are able to match the lower bound from above with an optimal test based on sparse eigenvalues. In the remaining region of the parameter space, where the dimensionality is high, the minimax lower bound implies that changepoint testing is very difficult. As our third contribution, we propose a computationally feasible variant of the optimal multivariate test for a change in covariance, which is also adaptive to the nominal noise level and the sparsity level of the change.
@@ 1 Introduction
Changepoint analysis, the problem of detecting or locating one or more distributional changes in a data sequence, has received renewed attention during the last decade. This surge in interest can be primarily attributed to its myriad applications in conjunction with the increasing availability of data. Distributional changes are of interest in several sciences as they can be interpreted as regime shifts. Examples include Economics (Andrews, 1993), Political Science (Djuve et al., 2020), Peace Research (Cunen et al., 2020) and Climate Research (Reeves et al., 2007), to mention a few. Another noteworthy application of changepoint analysis is condition monitoring, such as industrial (Pouliezos and Stavrakakis, 1994) or medical (Stival et al., 2023), where a changepoint may indicate a critical fault or abnormal behavior.
@@
Several theoretical contributions regarding the minimax properties of changepoint problems have been made in the past five years. Importantly, these contributions offer critical insight into the inherent difficulty of changepoint problems and serve as natural benchmarks for statistical methods. Most existing minimax results in the changepoint literature relate to changes in the mean vector. For instance, Liu et al. (2021) derive the exact minimax rate of testing for a single (possibly sparse) change in the mean vector with Gaussian noise, while Li et al. (2023) derive the minimax rate of testing in the more difficult scenario when the tails of the noise are either sub-Weibull or polynomially decaying. Other worthy mentions are the works of Enikeeva and Harchaoui (2019); Wang et al. (2020); Pilliat et al. (2023) and Verzelen et al. (2023), all focusing on changes in the mean vector.
@@
The story is somewhat different for problems involving variance or covariance changes. By a covariance change, we refer to a change in the spatial covariance matrix of the data, which for univariate data is simply a change in the variance. Such problems are fundamentally distinct from mean-change problems, as a covariance change alters the noise level of the data, unlike a change in the mean. To practitioners, covariance shifts are arguably of no less interest than mean shifts, as a covariance change entails a shift in either the variability of individual time series or their inter-dependencies. In finance, for example, the covariance between assets is of critical importance for portfolio construction (see, for instance, ŞErban et al. 2007 and Engle et al. 1990). In neuroscience, covariance and precision matrices are used quantify the functional connectivity between different regions of the brain, such as in Varoquaux et al. (2010); Cribben et al. (2013) and Jeong et al. (2016). Despite the large number of proposed methods for covariance changepoint detection, such as those of Aue et al. (2009); Galeano and Wied (2014) and Li and Li (2023), the fundamental difficulty of such problems have yet to be fully explored in the changepoint literature.
@@
Two exceptions are the recent works of Wang et al. (2021) and Enikeeva et al. (2023). The former derive a finite sample minimax lower bound on the signal strength at which no statistical method can reasonably estimate the location of a change in covariance in independent multivariate sub-Gaussian data. The lower bound does not cover the whole region of the parameter space, and does not capture the dependence on the sample size, nor any notion of sparsity. The latter paper considers a different setting with low rank VAR processes, and derive the asymptotic minimax testing rate for a change in the transition matrix of the observed process. In the sequel, we will only compare our results with those of Wang et al. (2021), as the setting investigated in Enikeeva et al. (2023) is too different from ours to obtain any meaningful comparisons. To best the author’s knowledge, no other contributions have been made to characterize the statistical hardness of changepoint problems with a covariance change.
@@
For stationary models, the inherent statistical hardness of variance and covariance related problems have been much more studied than for changepoint models. The works most related to ours are that of Berthet and Rigollet (2013) and Cai et al. (2015). The former propose a test for a sparse perturbation in the covariance matrix of i.i.d. Gaussian data based on $s$-sparse eigenvalues, and the latter show that this test is minimax rate optimal in the region of the parameter space where the signal strength is small or moderate. The current work builds upon the ideas and techniques in these two papers, although our problem setting, being a two sample problem, is fundamentally different than theirs. For other recent works, we refer to Cai et al. (2013), Cai et al. (2016) and Butucea and Zgheib (2016).
@@
In this paper, we investigate the minimax rate of testing for a change in the spatial covariance matrix of sequence of $n\geq 2$ independent, centered, sub-Gaussian observations of dimension $p$. We first consider the univariate case where $p=1$, in which case the spatial covariance matrix is simply the variance, under the assumption that the variance-rescaled data have bounded densities and the sub-Gaussian norm of the data is of the same order as the variance. Given a sequence of random variables $X_{1},\ldots,X_{n}$, we wish to test the null hypothesis that the sequence of variances $\mathbb{E}(X_{i}^{2})$ is constant versus the alternative that
@@
$\min(t_{0},n-t_{0})\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\min(\sigma_{1}^{2},\sigma_{2}^{2})}\right)\wedge\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\min(\sigma_{1}^{2},\sigma_{2}^{2})}\right)^{2}.$ (2)
Here, the symbol $\wedge$ indicates the minimum value. The first factor $\min(t_{0},n-t_{0})$ is the effective sample size of the problem, which is equal to the smallest of the number of observations before and after the change. The second factor measures the absolute size of the variance change relative to the smallest of the pre- and post-change variances.
@@
Next we consider multivariate data, assuming that the variance-rescaled data have bounded densities and the sub-Gaussian norm of the data is of the same order as the variance, now along any one-dimensional subspace of $\mathbb{R}^{p}$. Given a sequence $X_{1},\ldots,X_{n}$ of $p$-dimensional random variables, we wish to test the null hypothesis that $\mathbb{E}(X_{i}X_{i}^{\top})$ is constant and positive definite versus the alternative that
@@
where $S^{p-1}$ denotes the Euclidian unit sphere in $\mathbb{R}^{p}$. The right hand side of (5) is the largest absolute change in the variance of the data along some one-dimensional axis of variation. If an axis maximizing the change in variance is spanned by a vector with at most $s$ nonzero entries, we say that the change in covariance is $s$-sparse. That is, whenever the $s$-sparse eigenvalue
$\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})$ $=\underset{v\in S^{p-1}_{s}}{\sup}\ |v^{\top}(\Sigma_{1}-\Sigma_{2})v|$ (6)
agrees with the operator norm of $\Sigma_{1}-\Sigma_{2}$. Here, $S^{p-1}_{s}$ denotes the subspace of the Euclidian unit sphere containing only vectors with at most $s$ nonzero entries. As an example, the change in covariance is $s$-sparse whenever $\Sigma_{2}=\Sigma_{1}+vv^{\top}$, and $v$ has no more than $s$ nonzero entries. Similar notions of sparsity are used in e.g. Cai et al. (2015) and Berthet and Rigollet (2013).
@@
$\min(t_{0},n-t_{0})\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)\wedge\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)^{2}.$ (7)
@@
In a large region of the parameter space we match the minimax lower bound in (8) from above with an optimal test statistic. Specifically, this is the low- to moderate dimensional region where the effective sample size $\min(t_{0},n-t_{0})$ is no less than $C\left\{s\log(ep/s)\vee\log\log(8n)\right\}$, where the constant $C>0$ depends only on the desired testing level. In the high-dimensional region, where the effective sample size is even smaller, the lower bound still has interesting implications, even though the lower bound is not matched. In this region, the lower bound in (8) implies that the worst-case sum of Type I and Type II errors of any test can be made arbitrarily large whenever
$\underset{v\in\mathcal{S}^{p-1}_{s}}{\sup}\ \frac{v^{\top}\Sigma_{1}v}{v^{\top}\Sigma_{2}v}\vee\frac{v^{\top}\Sigma_{2}v}{v^{\top}\Sigma_{1}v}$ $\leq 1+c\frac{s\log\left(\frac{ep}{s}\right)\vee\log\log(8n)}{\min(t_{0},n-t_{0})},$ (9)
@@
by choosing the constant $c>0$ appropriately small. The left hand side of (9) is the maximal ratio of the pre- and post-change variances along any subspace of $\mathbb{R}^{p}$ spanned by an $s$-sparse vector. When $\min(t_{0},n-t_{0})$ is constant, the maximal ratio of the pre- and post-change variances must be of order $s\log(ep/s)\vee\log\log(8n)$ for changepoint detection to be possible, corresponding a stringent requirement on the signal strength of the changepoint.
@@
The optimal test statistic that we use to match the minimax lower bound in (8) is based on $s$-sparse eigenvalues of empirical covariance matrices, which are known to be NP-hard to compute. As the third contribution of this paper, we propose a computationally feasible, near-optimal test for a change in covariance that is adaptive to the sparsity and noise level of the data. The test is constructed via a convex relaxation of the sparse eigenvalue problem, which has a computational cost that is polynomial in $p$. This is done in Section 4.
@@
The proofs of our main results are given in Section 5, while auxiliary lemmas and proofs are given in the Supplementary Material. We use the following notation throughout the paper. For any set ${A}$, we let ${A}^{\complement}$ denote the complement of $A$ with respect to a universe $U\supseteq A$ determined from context. For any $d\in\mathbb{N}$ we let $[d]=\{1,2,\ldots,d\}$. For any pair $x,y\in\mathbb{R}$ we let $x\vee y=\max\{x,y\}$ and $x\wedge y=\min\{x,y\}$. Further, we let $\lfloor x\rfloor$ denote the largest integer no larger than $x$ and $\lceil x\rceil$ the smallest integer no smaller than $x$. We also write $x\lesssim y$ if $x\leq Cy$ for some context-dependent absolute constant $C>0$, and we write $x\asymp y$ if $y\lesssim x$ holds as well. For any vector $v=(v(1),\ldots,v(d))^{\top}\in\mathbb{R}^{d}$ and $p\geq 1$, we let $\left\lVert v\right\rVert_{p}$ denote the $\ell_{p}$-norm of $v$ given by $\left\lVert v\right\rVert_{p}=\{|v(1)|^{p}+\ldots+|v(d)|^{p}\}^{1/p}$. Moreover, we let $\left\lVert v\right\rVert_{\infty}=\max_{i\in[d]}\{v(i)\}$ denote the $\ell_{\infty}$ norm of $v$, and we let $\left\lVert v\right\rVert_{0}$ denote the $\ell_{0}$ ”norm” of $v$, which is given by number of nonzero elements in $v$. For any matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$ and $p\in\{0\}\cup[1,\infty]$, we let $\left\lVert A\right\rVert_{p}$ denote the $\ell_{p}$ norm of the vector formed by the entries of $A$. We define the operator norm of $A$ as $\left\lVert A\right\rVert_{\mathrm{op}}=\sup_{v\in\mathcal{S}^{d_{2}-1}}\left\lVert Ax\right\rVert_{2}$, where $\mathcal{S}^{d}$ denotes the Euclidian $d$-sphere. Moreover, if $A$ is symmetric, we define the largest absolute $s$-sparse eigenvalue of $A$ as $\lambda_{\max}^{s}(A)=\sup_{v\in S^{d_{2}-1}_{s}}|v^{\top}Av|$, where $S^{d}_{s}=\{v\in S^{d};\left\lVert v\right\rVert_{0}\leq s\}$ denotes the $s$-sparse Euclidian unit sphere. We also write $A\succcurlyeq 0$ to mean that $A$ is positive semi-definite. For any random variable $X$ taking values in $\mathbb{R}$, we let $\left\lVert X\right\rVert_{\Psi_{2}}=\inf\left\{t>0:\ \mathbb{E}\exp(X^{2}/t^{2})\leq 2\right\}$ denote the Orlicz-$\Psi_{2}$ norm of $X$, and we say that $X$ is sub-Gaussian if $\left\lVert X\right\rVert_{\Psi_{2}}<\infty$. For any $\mathbb{R}^{d}$-valued random variable $X$, we let $\left\lVert X\right\rVert_{\Psi_{2}}=\sup_{v\in\mathcal{S}^{d-1}}\left\lVert v^{\top}X\right\rVert_{\Psi_{2}}$ denote the Orlicz-$\Psi_{2}$ norm of $X$, saying that $X$ is sub-Gaussian if $\left\lVert X\right\rVert_{\Psi_{2}}<\infty$. For any two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on some measurable space $(\mathcal{X},\mathcal{A})$, we define the total variation distance to be $\text{TV}(\mathbb{P},\mathbb{Q})=\sup_{A\in\mathcal{A}}\ \left|\mathbb{P}(A)-\mathbb{Q}(A)\right|$. If $\mathbb{P}$ is absolutely continuous with respect to $\mathbb{Q}$, we define the Chi-square divergence to be $\chi^{2}(\mathbb{P}\ \|\ \mathbb{Q})=\int_{\mathcal{X}}(\mathrm{d}\mathbb{P}^{2}/\mathrm{d}\mathbb{Q}^{2})\mathrm{d}\mathbb{Q}-1$. Throughout we let $\mathbb{P}$ denote a generic probability measure determined from context, with associated expectation operator $\mathbb{E}$. When the probability measure is specified, say to be $P$, we let the expectation operator $\mathbb{E}_{P}$ and related functionals such as $\mathrm{Cov}_{P}(\cdot)$ be with respect to $P$.

### v2

@@ Abstract
We study the detection of a change in the covariance matrix of $n$ independent sub-Gaussian random variables of dimension $p$. Our first contribution is to show that $\log\log(8n)$ is the exact minimax testing rate for a change in variance when $p=1$, thereby giving a complete characterization of the problem for univariate data. Our second contribution is to derive a lower bound on the minimax testing rate under the operator norm, taking a certain notion of sparsity into account. In the low- to moderate-dimensional region of the parameter space, we are able to match the lower bound from above with an optimal test based on sparse eigenvalues. In the remaining region of the parameter space, where the dimensionality is high, the minimax lower bound implies that changepoint testing is very difficult. As our third contribution, we propose a computationally feasible variant of the optimal multivariate test for a change in covariance, which is also adaptive to the nominal noise level and the sparsity level of the change.
@@ 1 Introduction
Changepoint analysis, the problem of detecting or locating one or more distributional changes in a data sequence, has received renewed attention during the last decade. This surge in interest can be primarily attributed to its myriad applications in conjunction with the increasing availability of data. Distributional changes are of interest in several sciences as they can be interpreted as regime shifts. Examples include Economics (Andrews,, 1993), Political Science (Djuve et al.,, 2020), Peace Research (Cunen et al.,, 2020) and Climate Research (Reeves et al.,, 2007), to mention a few. Another noteworthy application of changepoint analysis is condition monitoring, such as industrial (Pouliezos and Stavrakakis,, 1994) or medical (Stival et al.,, 2023), where a changepoint may indicate a critical fault or abnormal behavior.
@@
Several theoretical contributions regarding the minimax properties of changepoint problems have been made in the past five years. Importantly, these contributions offer critical insight into the inherent difficulty of changepoint problems and serve as natural benchmarks for statistical methods. Most of the existing minimax results in the changepoint literature relate to changes in the mean vector. For instance, Liu et al., (2021) derive the exact minimax rate of testing for a single (possibly sparse) change in the mean vector with Gaussian noise, while Li et al., (2023) derive the minimax rate of testing in the more difficult scenario when the tails of the noise are either sub-Weibull or polynomially decaying. Other worthy mentions are the works of Enikeeva and Harchaoui, (2019); Wang et al., (2020); Pilliat et al., (2023) and Verzelen et al., (2023), all focusing on changes in the mean vector.
@@
The story is somewhat different for problems involving variance or covariance changes. By a covariance change, we refer to a change in the spatial covariance matrix of the data, which for univariate data is simply a change in the variance. Such problems are fundamentally distinct from mean-change problems, as a covariance change alters the noise level of the data, unlike a change in the mean. To practitioners, covariance shifts are arguably of no less interest than mean shifts, as a covariance change entails a shift in either the variability of individual time series or their inter-dependencies. In finance, for example, the covariance between assets is of critical importance for portfolio construction (see, for instance, ŞErban et al., 2007 and Engle et al., 1990). In neuroscience, covariance and precision matrices are used quantify the functional connectivity between different regions of the brain, such as in Varoquaux et al., (2010); Cribben et al., (2013) and Jeong et al., (2016). Despite the large number of proposed methods for covariance changepoint detection, such as those of Aue et al., (2009); Galeano and Wied, (2014) and Li and Li, (2023), the fundamental difficulty of such problems have yet to be fully explored in the changepoint literature.
@@
Two exceptions are the recent works of Wang et al., (2021) and Enikeeva et al., (2023). The former derive a finite sample minimax lower bound on the signal strength at which no statistical method can reasonably estimate the location of a change in covariance in independent multivariate sub-Gaussian data. The lower bound does not cover the whole region of the parameter space, and does not capture the dependence on the sample size, nor any notion of sparsity. The latter paper considers a different setting with low rank VAR processes, and derives the asymptotic minimax testing rate for a change in the transition matrix of the observed process. In the sequel, we will only compare our results with those of Wang et al., (2021), as the setting investigated in Enikeeva et al., (2023) is too different from ours to obtain any meaningful comparisons. To the best of the author’s knowledge, no other contributions have been made to characterize the statistical hardness of changepoint problems with a covariance change.
@@
For stationary models, the inherent statistical hardness of variance and covariance related problems have been much more studied than for changepoint models. The works most related to ours are that of Berthet and Rigollet, 2013b and Cai et al., (2015). The former propose a test for a sparse perturbation in the covariance matrix of i.i.d. Gaussian data based on $s$-sparse eigenvalues, and the latter show that this test is minimax rate optimal in the region of the parameter space where the signal strength is small or moderate. The current work builds upon the ideas and techniques in these two papers, although our problem setting, being a two-sample problem, is fundamentally different from theirs. For other works, we refer to Tony Cai and Xia, (2013), Cai et al., (2016) and Butucea and Zgheib, (2016).
@@
In this paper, we investigate the minimax rate of testing for a change in the covariance matrix of sequence of $n\geq 2$ independent, centered, sub-Gaussian observations of dimension $p$. We first consider the univariate case where $p=1$, in which case the covariance matrix is simply the variance, under the assumption that the variance-rescaled data have bounded densities and the sub-Gaussian norm of the data is of the same order as the variance. Given a sequence of random variables $X_{1},\ldots,X_{n}$, we wish to test the null hypothesis that the sequence of variances $\mathbb{E}(X_{i}^{2})$ is constant versus the alternative that
@@
$\min(t_{0},n-t_{0})\left\{\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\min(\sigma_{1}^{2},\sigma_{2}^{2})}\right)\wedge\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\min(\sigma_{1}^{2},\sigma_{2}^{2})}\right)^{2}\right\}.$ (2)
Here, the symbol $\wedge$ indicates the minimum value. The first factor $\min(t_{0},n-t_{0})$ is the effective sample size of the problem, which is equal to the smallest of the number of observations before and after the change. The second factor measures the absolute size of the variance change relative to the smallest of the pre- and post-change variances. We remark that $|\sigma_{1}^{2}-\sigma_{2}^{2}|/\min(\sigma_{1}^{2},\sigma_{2}^{2})$ can equivalently be written as the variance ratio $\{(\sigma_{1}^{2}/\sigma_{2}^{2})\vee(\sigma_{2}^{2}/\sigma_{1}^{2})\}-1$.
@@
Next we consider multivariate data, assuming that the variance-rescaled data have bounded densities and the sub-Gaussian norm of the data is of the same order as the variance, now along any one-dimensional subspace of $\mathbb{R}^{p}$. Given an independent sequence $X_{1},\ldots,X_{n}$ of $p$-dimensional random variables, we wish to test the null hypothesis that $\mathbb{E}(X_{i}X_{i}^{\top})$ is constant and positive definite versus the alternative that
@@
where $S^{p-1}$ denotes the Euclidean unit sphere in $\mathbb{R}^{p}$. The right-hand side of (5) is the largest absolute change in the variance of the data along some one-dimensional axis of variation. If an axis maximizing the change in variance is spanned by a vector with at most $s$ non-zero entries, we say that the change in covariance is $s$-sparse. That is, whenever the $s$-sparse eigenvalue $\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})$ agrees with the operator norm of $\Sigma_{1}-\Sigma_{2}$, defined by
$\lambda_{\max}^{s}(A)$ $=\underset{v\in S^{p-1}_{s}}{\sup}\ |v^{\top}Av|$ (6)
for any symmetric matrix $A$, where $S^{p-1}_{s}$ denotes the subspace of the Euclidean unit sphere containing only vectors with at most $s$ non-zero entries. As an example, the change in covariance is $s$-sparse whenever $\Sigma_{2}=\Sigma_{1}+vv^{\top}$, and $v$ has no more than $s$ non-zero entries. Similar notions of sparsity are used by e.g. Cai et al., (2015) and Berthet and Rigollet, 2013b for sparse PCA problems.
@@
$\min(t_{0},n-t_{0})\left\{\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)\wedge\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)^{2}\right\}.$ (7)
@@
In a large region of the parameter space we match the minimax lower bound in (8) from above with an optimal test statistic. Specifically, this is the low- to moderate-dimensional region where the effective sample size $\min(t_{0},n-t_{0})$ is no less than $C\left\{s\log(ep/s)\vee\log\log(8n)\right\}$, where the constant $C>0$ depends only on the desired testing level. In the high-dimensional region, where the effective sample size is even smaller, the lower bound still has interesting implications, even though the lower bound is not matched and thus not known to be tight. In this region, the lower bound in (8) implies that the worst-case sum of Type I and Type II errors of any test can be made arbitrarily large whenever
$\underset{v\in S^{p-1}_{s}}{\sup}\ \frac{v^{\top}\Sigma_{1}v}{v^{\top}\Sigma_{2}v}\vee\frac{v^{\top}\Sigma_{2}v}{v^{\top}\Sigma_{1}v}$ $\leq 1+c\frac{s\log\left(\frac{ep}{s}\right)\vee\log\log(8n)}{\min(t_{0},n-t_{0})},$ (9)
@@
by choosing the constant $c>0$ appropriately small. The left-hand side of (9) is the maximal ratio of the pre- and post-change variances along any subspace of $\mathbb{R}^{p}$ spanned by an $s$-sparse vector. When $\min(t_{0},n-t_{0})$ is constant, the maximal ratio of the pre- and post-change variances must be of order $s\log(ep/s)\vee\log\log(8n)$ for changepoint detection to be possible, corresponding a stringent requirement on the signal strength of the changepoint.
@@
The optimal test statistic that we use to match the minimax lower bound in (8) is based on $s$-sparse eigenvalues of empirical covariance matrices, which are known to be NP-hard to compute (Magdon-Ismail,, 2017). As the third contribution of this paper, we propose a computationally feasible, near-optimal test for a change in covariance that is adaptive to the sparsity and noise level of the data. The test is constructed via a convex relaxation of the sparse eigenvalue problem using Semidefinite Programming, which has a computational cost that is polynomial in $p$. This is done in Section 4.
@@
Proofs of all mathematical results in the paper, along with various auxiliary lemmas, are given in the Supplementary Material. We use the following notation throughout the paper. For any set ${A}$, we let ${A}^{\complement}$ denote the complement of $A$ with respect to a universe $U\supseteq A$ determined from context. For any $d\in\mathbb{N}$ we let $[d]=\{1,2,\ldots,d\}$. For any pair $x,y\in\mathbb{R}$ we let $x\vee y=\max\{x,y\}$ and $x\wedge y=\min\{x,y\}$. Further, we let $\lfloor x\rfloor$ denote the largest integer no larger than $x$ and $\lceil x\rceil$ the smallest integer no smaller than $x$. We also write $x\lesssim y$ if $x\leq Cy$ for some context-dependent absolute constant $C>0$, and we write $x\asymp y$ if $y\lesssim x$ holds as well. For any vector $v=(v(1),\ldots,v(d))^{\top}\in\mathbb{R}^{d}$ and $p\geq 1$, we let $\left\lVert v\right\rVert_{p}$ denote the $\ell_{p}$-norm of $v$ given by $\left\lVert v\right\rVert_{p}=\{|v(1)|^{p}+\ldots+|v(d)|^{p}\}^{1/p}$. Moreover, we let $\left\lVert v\right\rVert_{\infty}=\max_{i\in[d]}|v(i)|$ denote the $\ell_{\infty}$ norm of $v$, and we let $\left\lVert v\right\rVert_{0}$ denote the $\ell_{0}$ ”norm” of $v$, which is given by number of non-zero elements in $v$. For any matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$ and $p\in\{0\}\cup[1,\infty]$, we let $\left\lVert A\right\rVert_{p}$ denote the $\ell_{p}$ norm of the vector formed by the entries of $A$. We define the operator norm of $A$ as $\left\lVert A\right\rVert_{\mathrm{op}}=\sup_{v\in S^{d_{2}-1}}\left\lVert Ax\right\rVert_{2}$, where $S^{d}$ denotes the Euclidean $d$-sphere. Moreover, if $A$ is symmetric, we define the largest absolute $s$-sparse eigenvalue of $A$ as $\lambda_{\max}^{s}(A)=\sup_{v\in S^{d_{2}-1}_{s}}|v^{\top}Av|$, where $S^{d}_{s}=\{v\in S^{d};\left\lVert v\right\rVert_{0}\leq s\}$ denotes the $s$-sparse Euclidean unit sphere. We also write $A\succcurlyeq 0$ to mean that $A$ is positive semi-definite. For any random variable $X$ taking values in $\mathbb{R}$, we let $\left\lVert X\right\rVert_{\Psi_{2}}=\inf\left\{t>0:\ \mathbb{E}\exp(X^{2}/t^{2})\leq 2\right\}$ denote the Orlicz-$\Psi_{2}$ norm of $X$, and we say that $X$ is sub-Gaussian if $\left\lVert X\right\rVert_{\Psi_{2}}<\infty$. For any $\mathbb{R}^{d}$-valued random variable $X$, we let $\left\lVert X\right\rVert_{\Psi_{2}}=\sup_{v\in S^{d-1}}\left\lVert v^{\top}X\right\rVert_{\Psi_{2}}$ denote the Orlicz-$\Psi_{2}$ norm of $X$, saying that $X$ is sub-Gaussian if $\left\lVert X\right\rVert_{\Psi_{2}}<\infty$. For any two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on some measurable space $(\mathcal{X},\mathcal{A})$, we define the total variation distance to be $\text{TV}(\mathbb{P},\mathbb{Q})=\sup_{A\in\mathcal{A}}\ \left|\mathbb{P}(A)-\mathbb{Q}(A)\right|$. If $\mathbb{P}$ is absolutely continuous with respect to $\mathbb{Q}$, we define the Chi-square divergence to be $\chi^{2}(\mathbb{P}\ \|\ \mathbb{Q})=\int_{\mathcal{X}}(\mathrm{d}\mathbb{P}^{2}/\mathrm{d}\mathbb{Q}^{2})\mathrm{d}\mathbb{Q}-1$. Throughout we let $\mathbb{P}$ denote a generic probability measure determined from context, with associated expectation operator $\mathbb{E}$. When the probability measure is specified, say to be $P$, we let the expectation operator $\mathbb{E}_{P}$ and related functionals such as $\mathrm{Cov}_{P}(\cdot)$ be with respect to $P$.